OpenStudy (alivejeremy):

A. 6 cubic inches B. 8 cubic inches C. 12 cubic inches D. 16 cubic inches

OpenStudy (alivejeremy):

http://prntscr.com/a7rnxl

OpenStudy (alivejeremy):

Now u should remember the steps

OpenStudy (tamara7301):

B?

OpenStudy (alivejeremy):

So u tell me the steps now

OpenStudy (tamara7301):

um..

OpenStudy (alivejeremy):

B is right

OpenStudy (alivejeremy):

But tell me the steps

OpenStudy (tamara7301):

o ok..

OpenStudy (tamara7301):

find the length width and height

OpenStudy (tamara7301):

the multiply them

OpenStudy (alivejeremy):

Yeah

OpenStudy (alivejeremy):

length * width* height

OpenStudy (tamara7301):

4 *2=8 ^ 2*2*2=

OpenStudy (alivejeremy):

Yup

OpenStudy (alivejeremy):

Ur SMART :D

OpenStudy (tamara7301):

like that?

OpenStudy (tamara7301):

no im not -_-

OpenStudy (tamara7301):

u just taught me

OpenStudy (alivejeremy):

\[2\times2\times2\]

OpenStudy (alivejeremy):

\[2\times2\times =4\]

OpenStudy (tamara7301):

mhm

OpenStudy (alivejeremy):

\[2\times2=4\times2=8\]

OpenStudy (alivejeremy):

:D

OpenStudy (tamara7301):

smarty pants

OpenStudy (alivejeremy):

I medal u cuz u SMART

OpenStudy (tamara7301):

not

OpenStudy (alivejeremy):

u ARE

OpenStudy (alivejeremy):

i showed u one time u got it right the second time

OpenStudy (tamara7301):

not u just taught me remember

OpenStudy (tamara7301):

yea cuz u showed me ur the smart one not me

OpenStudy (alivejeremy):

We both r

OpenStudy (tamara7301):

lol sure

OpenStudy (alivejeremy):

lol

OpenStudy (tamara7301):

so wat now

OpenStudy (alivejeremy):

ok thts all for right now

OpenStudy (tamara7301):

o ok

OpenStudy (alivejeremy):

But u would get sum more

OpenStudy (tamara7301):

?

OpenStudy (alivejeremy):

More math silly

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

now pm me

OpenStudy (tamara7301):

o lol ok and y should i

OpenStudy (tamara7301):

;p

OpenStudy (alivejeremy):

Cuz this is spam if we talk here

OpenStudy (alivejeremy):

:D

OpenStudy (tamara7301):

o ok

OpenStudy (tamara7301):

then u pm me im to lazy;p